The particle displacement y of air molecules due to a sound wave is given by

y=0.008coswtsinkz where k=4π÷m and w=50πrads/s.

Calculate:

I) the distance between 2 consecutive nodes

ii) the amplitude after 0.565s



1
Expert's answer
2021-04-26T17:05:14-0400

i) the distance between 2 consecutive nodes is


"\lambda=\frac{2\pi}{k}=m."


ii) the amplitude after 0.565s is


"A_y=0.008\cos(50\pi·0.565)=1.75·10^{-4}\text{ m}."